OpenStudy (anonymous):

How can the alternative definition of power, \(P = \frac{w}{\Delta t} = F \frac{d}{\Delta t} \), can be derived by substituting the definitions of work and speed into the standard definition of power, \(P =\frac{W}{\Delta t}\).

OpenStudy (irishboy123):

power is the rate of change of work with time. that is P = dW/dt if the work is a force moving over a distance, then you can say \[dW = F dx,\space\ dP = \frac{d ( F dx)}{ dt} = \frac{dF}{dt} dx + F \frac{dx}{dt}\]. it all depends on what you are trying to achieve, really. you can measure energy conversion, say, the movement of a particle, as power. thus \[P = \frac{ d(\frac{1}{2} m v^{2} )} {dt} = \frac{1}{2} \frac{dm}{dt} v^{2} + mv \frac{dv}{dt} \] [although dm/dt might have an unusual meaning.]